OpenStudy (anonymous):

find the derivative of 10xsinx

OpenStudy (loser66):

product rule. DAT SIT

OpenStudy (anonymous):

hope it helps

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf \frac{d }{dx}(~~ 10x~sin(x) ~~)= }\) applying the product rule, \(\large\color{blue}{ \bf \frac{d }{dx} (~10x~)~~sin(x) +\frac{d }{dx} (~sin(x)~)~~10x }\) applying the power rule, \(\large\color{blue}{ \bf 10~~sin(x) +cos(x)~~10x }\) and there is the answer, \(\large\color{blue}{ \bf 10~~sin(x) +10~x~cos(x) }\) or in alternative form, \(\large\color{blue}{ \bf 10~(~sin(x) +~x~cos(x)~~) }\)